How many grams of calcium chloride were used if 2.442 grams of calcium oxalate monohydrate were produced, assuming a 100% yield. This means all of the calcium chloride was converted to calcium oxalate monohydrate. Report your answer to 3 decimal places in grams.
Answer #1

calcium oxalate monohydrate = CaC2O4·(H2O)

1 mole CaC2O4·(H2O) = 1 mol Ca

1 mole Ca = 1 mol CaCl2

1 mole CaC2O4·(H2O) = 1 mol CaCl2

No of mol of CaC2O4·(H2O) formed = w/Mwt = 2.442/146 = 0.0167 mole

No of mol of CaCl2 used = 0.0167 mole

mass of CaCl2 used = 0.0167*111 = 1.854 g